One bus traveles 20 km/hr faster than another. While the slower bus traveles 280 km the faster one goes 440 km. Find the speed of the faster bus.
a) 35 km/hr
b) 45 km/hr
c) 55 km/hr
d) 65 km/hr
e) none of these

Answer by Nate(3500)

I am guessing they traveled those distances in the same time (most likely scenerio.)
Distance = Rate * Time
Distance/Rate = Time
Times were equal:
Distance/Rate = Distance/Rate
Slower bus's distance: 280
Slower bus's rate: r
Faster bus's distance: 440
Faster bus's rate: r + 20
280/r = 440/(r + 20)
280(r + 20) = 440r
280r + 5600 = 440r
5600 = 160r
35 = r
Faster bus's rate: 35 + 20 = 55km/hr